OpenStudy (nomi):

A manufacturer believes that the net profit P is related to the advertising expenditure x by P(x)=10+46x-(1/2)x^2 what advertising expenditure will maximize the profit

OpenStudy (anonymous):

Take the derivative of P(x) \[P'(x)=-x+46\] To find the maximum, set the derivative equal to zero. \[0=-x+46\]\[x=46\]

OpenStudy (anonymous):

Are you in calculus? If not, there is also an algebraic way to do this.

OpenStudy (nomi):

algebra

OpenStudy (anonymous):

Oh, shoot, sorry. Alright, we know that the x-coordinate of the vertex is located at \[x=-\frac{b}{2a}\] So we need to figure out a, b, c before we can use that formula. Let's put it in the form \[0=ax^2+bx+c\]\[0=-\frac{1}{2}x+46x+10\]This means a=-1/2, b=46, c=10. \[x=-\frac{b}{2a}\]\[x=-\frac{46}{-(2)(\frac{1}{2})}=(-46)(-1)=46\]\[x=46\]

OpenStudy (nomi):

nice awesome so you change the expression to quadratic right?

OpenStudy (anonymous):

I put it in general quadratic form. It was a quadratic before, but the general form makes it easy to identify what a, b, and c are.